# PSDNorm: Test-Time Temporal Normalization for Deep Learning in Sleep Staging

### v2

@@ 3.2 Implementation details
1:  Input: Batch =$\mathcal{B}=\left\{\mathbf{G}^{(1)},\dots,\mathbf{G}^{(N)}\right\}$, running barycenter $\widehat{\overline{\mathbf{P}}}$, filter-size $F$, momentum $\alpha$, training flag
@@
The forward computation of the proposed layer is outlined in Algorithm 1. At train time, the PSDNorm performs three main operations: 1) PSD estimation, 2) running Riemannian barycenter update, and 3) Monge mapping application. At test time, the PSDNorm operates similarly, except it does not update the running barycenter. The PSDNorm is fully differentiable and can be integrated into any deep learning model. Similarly to classical normalization layers, a stop gradient operation is applied to the running barycenter to prevent the backpropagation of the gradient computation through the barycenter. PSDNorm has a unique additional hyperparameter $F$ which is the filter size. It controls the alignment between each feature map and the running barycenter PSD and it is typically chosen in our experiments between 1 and 17. In practice, the Fourier transforms are efficiently computed using the Fast Fourier Transform (FFT) algorithm. Because of the estimation of PSDs, the complexity of the PSDNorm, both at train and test times, is $\mathcal{O}(Nc\ell F\log(F))$, where $N$ is the batch size, $c$ the number of channels, $\ell$ the signal length, and $F$ the filter size.
@@
InstanceNorm applies a per-channel $z$-score over time, subtracting the mean and dividing by the standard deviation—equivalent to whitening under an i.i.d. assumption over time. In contrast, PSDNorm explicitly accounts for temporal structure by estimating the PSD and whitening/re-coloring in the frequency domain. InstanceNorm is recovered as a special case of PSDNorm by setting the filter size to $F=1$ and using the identity as the re-coloring transform instead of the barycentric PSD.
@@ 4 Numerical Experiments
In this section, we evaluate the proposed method through a series of experiments designed to highlight its effectiveness and robustness on the clinically relevant task of sleep staging. We first describe the datasets and training setup employed, followed by a performance comparison with existing normalization techniques. Next, we assess the efficiency of PSDNorm by training over varying numbers of subjects per dataset. Finally, we analyze the robustness of PSDNorm against domain shift by focusing on subject-wise performance and different architectures. The code will be available on GitHub upon acceptance. The anonymized code is available in the supplementary material. All numerical experiments were conducted using a total of 1500 GPU hours on NVIDIA H100 GPUs.
@@ 4.1 Experimental Setup
Datasets
To evaluate the effect of normalization layers, we use ten datasets of sleep staging described in Table 1. ABC [33], CCSHS [34], CFS [35], HPAP [36], MROS [37], SHHS [4], CHAT [5], and SOF [38] are publicly available sleep datasets with restricted access from National Sleep Research Resource (NSRR) [39]. PHYS [40] and MASS [3] are two other datasets publicly available. Every 30 s epoch is labeled with one of the five sleep stages: Wake, N1, N2, N3, and REM. These datasets are unbalanced in terms of age, sex, number of subjects, and have been recorded with different sensors in different institutions which makes the sleep staging task challenging. We now describe the pre-processing steps and splits of the datasets.
@@
We follow a standard pre-processing pipeline used in the field [41, 42]. The datasets vary in the number and type of available EEG and electrooculogram (EOG) channels. To ensure consistency, we use two bipolar EEG channels, as some datasets lack additional channels. For dataset from NSRR, we select the channels C3-A2 and C4-A1. For signals from Physionet and MASS, we use the only available channels Fpz-Cz and Pz-Oz. The EEG signals are low-pass filtered with a 30 Hz cutoff frequency and resampled to 100 Hz. All data extraction and pre-processing steps are implemented using MNE-BIDS [43] and MNE-Python [44].
@@
Table 2: Balanced Accuracy (BACC) scores on the left-out datasets. The top section reports results in the large-scale setting (using all available subjects), while the bottom section presents results in the medium-scale setting (balanced@400). For each row, the best score is highlighted in bold, and standard deviations reflect training variability across 3 random seeds.
BatchNorm LayerNorm InstanceNorm PSDNorm(F=5) PSDNorm(F=9) PSDNorm(F=17)
All subjects ABC 78.49 $\pm$ 0.42 77.94 $\pm$ 0.31 78.83 $\pm$ 0.59 78.56 $\pm$ 0.67 78.73 $\pm$ 0.32 78.60 $\pm$ 0.28
CCSHS 88.79 $\pm$ 0.21 87.51 $\pm$ 0.77 88.75 $\pm$ 0.04 88.56 $\pm$ 0.36 88.48 $\pm$ 0.07 88.52 $\pm$ 0.19
CFS 84.97 $\pm$ 0.37 84.29 $\pm$ 0.67 85.73 $\pm$ 0.29 85.42 $\pm$ 0.09 85.25 $\pm$ 0.19 85.28 $\pm$ 0.12
CHAT 64.72 $\pm$ 3.94 64.36 $\pm$ 0.40 68.86 $\pm$ 2.49 70.57 $\pm$ 1.24 70.36 $\pm$ 2.22 70.71 $\pm$ 2.15
HOMEPAP 76.39 $\pm$ 0.29 75.23 $\pm$ 0.78 76.70 $\pm$ 0.35 76.72 $\pm$ 0.27 76.93 $\pm$ 0.10 77.02 $\pm$ 0.32
MASS 73.71 $\pm$ 0.62 71.39 $\pm$ 3.00 72.12 $\pm$ 0.70 72.51 $\pm$ 1.68 71.34 $\pm$ 2.68 71.61 $\pm$ 0.71
MROS 81.30 $\pm$ 0.25 80.44 $\pm$ 0.29 81.49 $\pm$ 0.18 81.57 $\pm$ 0.34 81.35 $\pm$ 0.17 81.30 $\pm$ 0.13
PhysioNet 76.13 $\pm$ 0.57 75.12 $\pm$ 0.22 76.15 $\pm$ 0.52 75.96 $\pm$ 1.02 76.35 $\pm$ 0.27 76.02 $\pm$ 0.47
SHHS 77.97 $\pm$ 1.46 75.98 $\pm$ 0.48 79.05 $\pm$ 0.89 79.14 $\pm$ 1.01 79.33 $\pm$ 0.62 79.12 $\pm$ 0.11
SOF 81.33 $\pm$ 0.54 81.82 $\pm$ 0.79 81.98 $\pm$ 0.22 82.50 $\pm$ 0.34 82.15 $\pm$ 1.00 81.94 $\pm$ 0.65
Mean 78.38 $\pm$ 0.47 77.41 $\pm$ 0.28 78.97 $\pm$ 0.11 79.15 $\pm$ 0.14 79.03 $\pm$ 0.10 79.01 $\pm$ 0.36
Balanced@400 ABC 78.26 $\pm$ 1.33 75.29 $\pm$ 0.81 78.73 $\pm$ 0.42 78.18 $\pm$ 0.68 78.18 $\pm$ 0.91 77.76 $\pm$ 1.00
CCSHS 87.42 $\pm$ 0.16 85.20 $\pm$ 0.48 87.62 $\pm$ 0.42 87.58 $\pm$ 0.30 87.35 $\pm$ 0.52 87.62 $\pm$ 0.48
CFS 84.32 $\pm$ 0.57 81.66 $\pm$ 1.36 84.72 $\pm$ 0.33 84.29 $\pm$ 0.36 84.06 $\pm$ 0.10 84.46 $\pm$ 0.06
CHAT 66.55 $\pm$ 0.88 61.19 $\pm$ 1.16 64.43 $\pm$ 4.41 70.28 $\pm$ 1.70 68.11 $\pm$ 3.94 69.88 $\pm$ 0.46
HOMEPAP 75.25 $\pm$ 0.50 74.86 $\pm$ 0.25 76.47 $\pm$ 0.63 76.83 $\pm$ 0.61 76.61 $\pm$ 0.74 76.49 $\pm$ 0.45
MASS 70.00 $\pm$ 1.91 68.56 $\pm$ 3.33 71.52 $\pm$ 1.13 72.77 $\pm$ 1.09 73.07 $\pm$ 1.30 72.23 $\pm$ 2.40
MROS 80.37 $\pm$ 0.20 78.05 $\pm$ 0.22 80.28 $\pm$ 0.21 80.26 $\pm$ 0.11 80.32 $\pm$ 0.22 80.70 $\pm$ 0.42
PhysioNet 75.81 $\pm$ 0.13 71.82 $\pm$ 2.12 74.68 $\pm$ 0.55 74.82 $\pm$ 2.11 73.77 $\pm$ 1.73 75.09 $\pm$ 0.97
SHHS 76.44 $\pm$ 0.92 75.12 $\pm$ 0.39 78.68 $\pm$ 0.37 78.88 $\pm$ 0.68 77.28 $\pm$ 0.91 78.41 $\pm$ 0.49
SOF 81.08 $\pm$ 1.14 78.70 $\pm$ 0.50 80.68 $\pm$ 1.38 79.49 $\pm$ 0.41 81.44 $\pm$ 0.97 81.07 $\pm$ 0.66
Mean 77.55 $\pm$ 0.34 75.05 $\pm$ 0.28 77.78 $\pm$ 0.46 78.34 $\pm$ 0.42 78.02 $\pm$ 0.67 78.37 $\pm$ 0.38
@@
Sleep staging has inspired a variety of neural architectures, from early CNN-based models [41, 42, 19] to recent attention-based approaches [21, 17, 20]. We evaluate two architectures: U-Sleep [45, 15], a state-of-the-art temporal CNN model designed for robustness and large-scale training, and a newly introduced architecture, CNNTransformer. CNNTransformer combines a lightweight convolutional encoder with a Transformer applied to epoch-level embeddings. It is specifically tailored for two-channel EEG and designed to scale efficiently to large datasets, while remaining minimal in implementation (under 100 lines of code) and training cost (Appendix A). Its design draws inspiration from recent transformer-based models for time series [46], with an emphasis on simplicity and practicality.
We use the Adam optimizer [47] with a learning rate of $10^{-3}$ to minimize the weighted cross-entropy loss, where class weights are computed from the training set distribution. Training is performed with a batch size of 64, and early stopping is applied based on validation loss with a patience of 3 epochs. Each input corresponds to a sequence of 17’30s, with a stride of 10’30s between sequences along the full-night recording.
@@
At test time, the model similarly processes sequences of 17’30s with a stride of 10’30s. Performance is evaluated using the balanced accuracy score (BACC), computed on the central 10’30s of each prediction window. Each experiment is repeated three times with different random seeds, and we report the mean and standard deviation of BACC.
@@
We compare the proposed PSDNorm with three normalization strategies: BatchNorm, LayerNorm, and InstanceNorm. Note that InstanceNorm corresponds to a special case of PSDNorm with $F=1$ and a fixed identity mapping instead of a learned running barycenter. In the following experiments, the BatchNorm layers in the first three convolutional layers are replaced with either PyTorch’s default implementations of LayerNorm, InstanceNorm [48], or PSDNorm. To preserve the receptive field, the filter size $F$ of PSDNorm is used in the first layer and progressively halved in the following ones. We fix the momentum $\alpha$ to $10^{-2}$.
@@ 4.2 Numerical Results
Table 2 (top) reports the LODO BACC of U-Sleep across all datasets, averaged over three random seeds. PSDNorm consistently outperforms all baseline normalization layers—BatchNorm, LayerNorm, and InstanceNorm—achieving the highest mean BACC of 79.15%, which exceeds BatchNorm (78.38%), InstanceNorm (78.97%), and LayerNorm (77.41%) by at least one standard deviation. This performance holds across tested filter sizes, with PSDNorm reaching 79.03% for $F{=}9$ and 79.01% for $F{=}17$. In total, PSDNorm ranks first on 6 out of the 10 datasets. On the challenging CHAT dataset, where all methods struggle, PSDNorm with $F{=}5$ outperforms all other normalizations by more than 2 percentage points, highlighting its robustness under strong distribution shifts. Although InstanceNorm is a strong baseline—outperforming BatchNorm and LayerNorm by at least one standard deviation on average—it is consistently surpassed by PSDNorm in average performance. In contrast, LayerNorm underperforms across the board, achieving the lowest average BACC and never ranking first, confirming its limited suitability for this task.
Efficiency: Performance with 4$\times$ Less Data
@@
The PSDNorm layer improves model performance when trained on the full dataset ($\sim$10000 subjects), but such large-scale data availability is not always the case. In many real-world scenarios—such as rare disease studies, pediatric populations, or data collected in constrained clinical settings—labeled recordings are scarce, expensive to annotate, or restricted due to privacy concerns. Evaluating model robustness under these constraints is therefore essential. To this end, we train all models using the balanced@400 setup, which reduces the training data by a factor of 4 compared to the full-data setting. In this lower-data regime, PSDNorm continues to outperform all baseline normalization strategies on 6 out of 10 datasets and achieves higher average BACC across all tested values of $F$. The performance improvement of PSDNorm over the best baseline is more pronounced in this setting: for $F{=}5$, the BACC gain reaches $+0.56\%$, compared to $+0.18\%$ in the full-data setting. Again, the gains exceed one standard deviation. To assess statistical significance, we conducted a critical difference (CD) test [49]. Figure 3 (top) reports the average rank of each method and the corresponding statistical comparisons. The results confirm that PSDNorm significantly outperforms the baselines, underscoring the value of incorporating temporal structure into normalization for robust and data-efficient generalization. The same trend is observed for U-Sleep trained on all subjects (see in Appendix Figure 6). The following experiments focus on the balanced@400 setup.
@@
PSDNorm is a plug-and-play normalization layer that can be seamlessly integrated into various neural network architectures. To demonstrate this flexibility, we evaluate its performance on both the U-Sleep and CNNTransformer models. Figure 3 reports the average rank of each normalization method across datasets and subjects for both architectures using datasets balanced@400. In both architectures, PSDNorm with $F{=}5$ achieves the best overall ranking and demonstrates statistically significant improvements over both BatchNorm, LayerNorm and InstanceNorm. Notably, it is also the only method to rank among the top two across both architectures. The results confirm that PSDNorm generalizes well beyond a single architecture and can provide consistent improvements in diverse modeling setups. InstanceNorm performs competitively in some cases but is never significantly better than PSDNorm. The figure also illustrates that PSDNorm with $F{=}17$ does not yield additional benefit in this medium-scale setting, further supporting the conclusion that moderate temporal context is sufficient at this data scale. Detailed numerical scores for CNNTransformer are reported in the supplementary material (Table 7).
Sensitivity to Filter Size
The choice of $F$ in PSDNorm controls the intensity of the normalization: larger $F$ provide stronger normalization, while smaller $F$ allow more flexibility in the model. In Figure 4, we evaluate its impact across different training set sizes and observe a clear trend: when trained on fewer subjects, larger filter sizes yield better performance (i.e., $F=17$), whereas smaller filter sizes are more effective with larger datasets (i.e., $F={5}$). This suggests that with limited data, stronger normalization helps prevent overfitting, while with more data, a more flexible model is preferred. On average, PSDNorm with $F=5$ offers a good compromise, achieving one of the best performances across all training set sizes.
@@
Performance variability across subjects is a key challenge in biomedical applications where ensuring consistently high performance—even for the most challenging subjects—is critical. To highlight the robustness of PSDNorm, Figure 5 presents a scatter plot of subject-wise BACC scores comparing BatchNorm or InstanceNorm vs. PSDNorm across two selected target datasets. CHAT and MASS are two challenging datasets, where the prediction performance is significantly lower than the other datasets. For CHAT, most of the dots are below the diagonal, indicating that PSDNorm improves performance for 91% of subjects against BatchNorm and 99% of subjects against InstanceNorm, with the largest gains observed for the hardest subjects, reinforcing its ability to handle challenging cases. For MASS, PSDNorm improves performance for 75% of subjects against BatchNorm and 69% against InstanceNorm. This demonstrates that PSDNorm is not only effective in improving overall performance but also excels in enhancing the performance of the most challenging subjects.

### v3

@@ 3.2 Implementation details
The forward computation of the proposed layer is outlined in Algorithm˜1. At train time, the PSDNorm performs three main operations: 1) PSD estimation, 2) running Riemannian barycenter update, and 3) Monge mapping application. At inference, the PSDNorm operates similarly, except it does not update the running barycenter. The PSDNorm is fully differentiable and can be integrated into any deep learning model. Similarly to classical normalization layers, a stop gradient operation is applied to the running barycenter to prevent the backpropagation of the gradient computation through the barycenter. PSDNorm has a unique additional hyperparameter $f$ which is the filter size. It controls the alignment between each feature map and the running barycenter PSD and it is typically chosen in our experiments between 1 and 17. In practice, the Fourier transforms are efficiently computed using the Fast Fourier Transform (FFT) algorithm. Because of the estimation of PSDs, the complexity of the PSDNorm, both at train and inference times, is $\mathcal{O}(Nc\ell f\log(f))$, where $N$ is the batch size, $c$ the number of channels, $\ell$ the signal length, and $f$ the filter size.
3.3 Discussion and Connections to Related Methods
@@
1:Input: Batch $\mathcal{B}=\left\{\mathbf{G}^{(1)},\dots,\mathbf{G}^{(N)}\right\}$, running barycenter $\widehat{\overline{\mathbf{P}}}$, filter-size $f$, momentum $\alpha$, training flag
@@
InstanceNorm applies a per-channel $z$-score over time, subtracting the mean and dividing by the standard deviation—equivalent to whitening under an i.i.d. assumption over time. In contrast, PSDNorm explicitly accounts for temporal structure by estimating the PSD and whitening/re-coloring in the frequency domain. InstanceNorm is recovered as a special case of PSDNorm by setting the filter size to $f=1$ and using the uniform PSD barycenter as $\widehat{\overline{\mathbf{P}}}={\mathbf{1}}$. as the re-coloring transform instead of the barycentric PSD.
Similarity with Test-time Domain Adaptation
PSDNorm is inspired by Temporal Monge Alignment (TMA) (Gnassounou et al., 2023), a pre-processing technique that can be used for test-time adaptation. Test-time Domain Adaptation methods adjust a pre-trained model to a new target domain during inference, without requiring access to the original training data (Wang et al., 2021; Yang et al., 2021). While PSDNorm must be integrated into the model during training and is not a post-hoc adaptation method, it provides a similar benefit at inference time. Designing new modern architectures that incorporate PSDNorm can enhance robustness to domain shifts without the need for retraining or access to source data.
Discussion of Gaussian and Stationarity Assumptions
PSDNorm relies on the Gaussian approximation of OT for compensation variability but does not assume that the signals are Gaussian. This allows for efficient alignment of second-order statistics (covariance structure), but also allows preserving higher-order discriminative information. This approach is computationally tractable and targets the most prominent sources of domain shift without over-distorting the signal, a strategy also used in successful methods like Deep CORAL (Sun and Saenko, 2016). Like BatchNorm, PSDNorm assumes shifts are captured by low-order statistics, but it provides a richer alignment by incorporating temporal context.
@@ 4 Numerical Experiments
In this section, we evaluate the proposed method through a series of experiments designed to highlight its effectiveness and robustness on the clinically relevant task of sleep staging. We first describe the datasets and training setup employed, followed by a performance comparison with existing normalization techniques. Next, we assess the efficiency of PSDNorm by training over varying numbers of subjects per dataset. Finally, we analyze the robustness of PSDNorm against domain shift by focusing on subject-wise performance and different architectures. The code is available at https://github.com/tgnassou/PSDNorm. The anonymized code is available in the supplementary material. All numerical experiments were conducted using a total of 1500 GPU hours on NVIDIA H100 GPUs.
@@ 4.1 Experimental Setup
To evaluate the effect of normalization layers, we use ten datasets of sleep staging described in Table˜1. ABC (Jessie P. et al., 2018), CCSHS (Rosen et al., 2003), CFS (Redline et al., 1995), HPAP (Rosen et al., 2012), MROS (Blackwell et al., 2011), SHHS (Quan et al., 1998), CHAT (Marcus et al., 2013), and SOF (Spira et al., 2008) are publicly available sleep datasets with restricted access from National Sleep Research Resource (NSRR) (Zhang et al., 2018). PHYS (Goldberger et al., 2000) and MASS (O’Reilly et al., 2014) are two other datasets publicly available. Every 30 s epoch is labeled with one of the five sleep stages: Wake, N1, N2, N3, and REM. These datasets are unbalanced in terms of age, sex, number of subjects, and have been recorded with different sensors in different institutions which makes the sleep staging task challenging. We now describe the pre-processing steps and splits of the datasets.
@@
We follow a standard pre-processing pipeline used in the field (Chambon et al., 2017; Stephansen et al., 2018). The datasets vary in the number and type of available EEG and electrooculogram (EOG) channels. To ensure consistency, we use two bipolar EEG channels, as some datasets lack additional channels. For dataset from NSRR, we select the channels C3-A2 and C4-A1. For signals from Physionet and MASS, we use the only available channels Fpz-Cz and Pz-Oz. The EEG signals are low-pass filtered with a 30 Hz cutoff frequency and resampled to 100 Hz. All data extraction and pre-processing steps are implemented using MNE-BIDS (Appelhoff et al., 2019) and MNE-Python (Gramfort et al., 2013).
@@
Table 2: Balanced Accuracy (BACC) scores on the left-out datasets with USleep. The top section reports results in the large-scale setting (using all available subjects), while the bottom section presents results in the medium-scale setting (balanced@400). For each row, the best score is highlighted in bold, and standard deviations reflect training variability across 3 random seeds. The mean BACC reports the average over all the subjects.
Dataset BatchNorm LayerNorm InstanceNorm TMA PSDNorm
All subjects ABC $8.49_{\pm 0.42} 77$$7.94_{\pm 0.31}$ $\mathbf{78.83_{\pm 0.59}}$ $78.33_{\pm 0.12}$ $78.56_{\pm 0.67}$
CCSHS $\mathbf{88.79_{\pm 0.21}}$ $7.51_{\pm 0.77} 88$$75_{\pm 0.04} 88.$$61_{\pm 0.10} 88.$$8.56_{\pm 0.36}$
CFS $84.97_{\pm 0.37}$ $84.29_{\pm 0.67}$ $\mathbf{85.73_{\pm 0.29}}$ $84.85_{\pm 0.13}$ $85.42_{\pm 0.09}$
CHAT $64.72_{\pm 3.94}$ $64.36_{\pm 0.40}$ $68.86_{\pm 2.49}$ $69.76_{\pm 1.62}$ $\mathbf{70.57_{\pm 1.24}}$
HOMEPAP $76.39_{\pm 0.29}$ $75.23_{\pm 0.78}$ $76.70_{\pm 0.35}$ $\mathbf{76.77_{\pm 0.66}}$ $76.72_{\pm 0.27}$
MASS $73.71_{\pm 0.62}$ $71.39_{\pm 3.00}$ $72.12_{\pm 0.70}$ $\mathbf{73.90_{\pm 0.69}}$ $72.51_{\pm 1.68}$
MROS $81.30_{\pm 0.25}$ $80.44_{\pm 0.29}$ $81.49_{\pm 0.18}$ $80.91_{\pm 0.42}$ $\mathbf{81.57_{\pm 0.34}}$
PhysioNet $76.13_{\pm 0.57}$ $75.12_{\pm 0.22}$ $76.15_{\pm 0.52}$ $\mathbf{76.48_{\pm 0.37}}$ $75.96_{\pm 1.02}$
SHHS $77.97_{\pm 1.46}$ $75.98_{\pm 0.48}$ $79.05_{\pm 0.89}$ $78.21_{\pm 0.39}$ $\mathbf{79.14_{\pm 1.01}}$
SOF $81.33_{\pm 0.54}$ $81.82_{\pm 0.79}$ $81.98_{\pm 0.22}$ $81.84_{\pm 0.49}$ $\mathbf{82.50_{\pm 0.34}}$
Mean(Dataset) $8.38_{\pm 0.47} 77$$7.41_{\pm 0.28}$ $78.97_{\pm 0.11}$ $78.98_{\pm 0.14}$ $\mathbf{79.15_{\pm 0.14}}$
Mean(Subject) $78.14_{\pm 1.01}$ $76.78_{\pm 0.18}$ $79.26_{\pm 0.48}$ $78.77_{\pm 0.07}$ $\mathbf{79.51_{\pm 0.62}}$
Balanced@400 ABC $78.26_{\pm 1.33}$ $75.29_{\pm 0.81}$ $\mathbf{78.73_{\pm 0.42}}$ $78.04_{\pm 0.51}$ $78.18_{\pm 0.68}$
CCSHS $87.42_{\pm 0.16}$ $85.20_{\pm 0.48}$ $\mathbf{87.62_{\pm 0.42}}$ $87.57_{\pm 0.20}$ $87.58_{\pm 0.30}$
CFS $84.32_{\pm 0.57}$ $81.66_{\pm 1.36}$ $\mathbf{84.72_{\pm 0.33}}$ $84.58_{\pm 0.20}$ $84.29_{\pm 0.36}$
CHAT $66.55_{\pm 0.88}$ $61.19_{\pm 1.16}$ $64.43_{\pm 4.41}$ $68.73_{\pm 2.48}$ $\mathbf{70.28_{\pm 1.70}}$
HOMEPAP $75.25_{\pm 0.50}$ $74.86_{\pm 0.25}$ $76.47_{\pm 0.63}$ $76.10_{\pm 0.32}$ $\mathbf{76.83_{\pm 0.61}}$
MASS $70.00_{\pm 1.91}$ $68.56_{\pm 3.33}$ $71.52_{\pm 1.13}$ $71.63_{\pm 1.92}$ $\mathbf{72.77_{\pm 1.09}}$
MROS $\mathbf{80.37_{\pm 0.20}}$ $78.05_{\pm 0.22}$ $80.28_{\pm 0.21}$ $80.09_{\pm 0.40}$ $80.26_{\pm 0.11}$
PhysioNet $\mathbf{75.81_{\pm 0.13}}$ $71.82_{\pm 2.12}$ $74.68_{\pm 0.55}$ $75.31_{\pm 1.54}$ $74.82_{\pm 2.11}$
SHHS $76.44_{\pm 0.92}$ $75.12_{\pm 0.39}$ $8.68_{\pm 0.37} 77$$7.00_{\pm 0.39}$ $\mathbf{78.88_{\pm 0.68}}$
SOF $81.08_{\pm 1.14}$ $78.70_{\pm 0.50}$ $80.68_{\pm 1.38}$ $\mathbf{81.25_{\pm 0.71}}$ $79.49_{\pm 0.41}$
Mean(Dataset) $77.55_{\pm 0.34}$ $5.05_{\pm 0.28} 77$$7.78_{\pm 0.46}$ $78.03_{\pm 0.35}$ $\mathbf{78.34_{\pm 0.42}}$
Mean(Subject) $77.22_{\pm 0.34}$ $75.04_{\pm 0.42}$ $8.17_{\pm 0.28} 77$$7.74_{\pm 0.36}$ $\mathbf{78.85_{\pm 0.59}}$
@@
Sleep staging has inspired a variety of neural architectures, from early CNN-based models (Chambon et al., 2017; Stephansen et al., 2018; Phan et al., 2022a) to recent attention-based approaches (Phan et al., 2022b; 2023; Wang et al., 2024). We evaluate two architectures: U-Sleep (Perslev et al., 2019; 2021), a state-of-the-art temporal CNN model designed for robustness and large-scale training, and a newly introduced architecture, CNNTransformer. CNNTransformer combines a lightweight convolutional encoder with a Transformer applied to epoch-level embeddings. It is specifically tailored for two-channel EEG and designed to scale efficiently to large datasets, while remaining minimal in implementation (under 100 lines of code) and training cost (Section˜A.4). Its design draws inspiration from recent transformer-based models for time series (Yang et al., 2023), with an emphasis on simplicity and practicality.
We use the Adam optimizer (Kingma, 2014) with a learning rate of $10^{-3}$ to minimize the weighted cross-entropy loss, where class weights are computed from the training set distribution. Training is performed with a batch size of 64, and early stopping is applied based on validation loss with a patience of 3 epochs. Each input corresponds to a sequence of 17’30s, with a stride of 10’30s between sequences along the full-night recording. The filter size $f$ of PSDNorm is set to 5. A sensitivity analysis of $f$ is provided in Section˜A.6 in the appendix, and show that the performance is stable across a range of values from 5 to 11.
@@
At inference, the model similarly processes sequences of 17’30s with a stride of 10’30s. Performance is evaluated using the balanced accuracy score (BACC), computed on the central 10’30s of each prediction window. Each experiment is repeated three times with different random seeds, and we report the mean and standard deviation of BACC.
@@
We compare the proposed PSDNorm with three normalization strategies: BatchNorm, LayerNorm, and InstanceNorm. Note that InstanceNorm corresponds to a special case of PSDNorm with $f=1$ and a fixed identity mapping instead of a learned running barycenter. In the following experiments, the BatchNorm layers in the first three convolutional layers are replaced with either PyTorch’s default implementations of LayerNorm, InstanceNorm (Paszke et al., 2019), or PSDNorm. To preserve the receptive field, the filter size $f$ of PSDNorm is used in the first layer and progressively halved in the following ones. We fix the momentum $\alpha$ to $10^{-2}$.
@@ 4.2 Numerical Results
Table˜2 (top) reports the LODO BACC of U-Sleep across all datasets, averaged over three random seeds. PSDNorm consistently outperforms all baseline normalization layers—BatchNorm, LayerNorm, InstanceNorm, and TMA—achieving the highest mean BACC of 79.51% over subjects, which exceeds BatchNorm (78.38%), InstanceNorm (78.97%), LayerNorm (77.41%) and TMA (78.77%). On the challenging CHAT dataset, where all methods struggle, PSDNorm outperforms all other normalizations by more than 1 percentage points, highlighting its robustness under strong distribution shifts. Although InstanceNorm is a strong baseline—outperforming BatchNorm and LayerNorm by at least one standard deviation on average—it is consistently surpassed by PSDNorm in average performance. In contrast, LayerNorm underperforms across the board, achieving the lowest average BACC and never ranking first, confirming its limited suitability for this task. PSDNorm also improves score by almost 1% over TMA, showing that using Monge Alignment inside the network allows for better adaptation.
[…]
@@
The PSDNorm layer improves model performance when trained on the full dataset ($\sim$10000 subjects), but such large-scale data availability is not always the case. In many real-world scenarios—such as rare disease studies, pediatric populations, or data collected in constrained clinical settings—labeled recordings are scarce, expensive to annotate, or restricted due to privacy concerns. Evaluating model robustness under these constraints is therefore essential. To this end, we train all models using the balanced@400 setup, which reduces the training data by a factor of 4 compared to the full-data setting. In this lower-data regime, PSDNorm continues to outperform all baseline normalization strategies and achieves higher average BACC. The performance improvement of PSDNorm over the best baseline is more pronounced in this setting: the BACC gain reaches $0.67\%$, compared to $0.25\%$ in the full-data setting. The gains exceed one standard deviation. To assess statistical significance, we conducted a critical difference (CD) test (Demšar, 2006). Figure˜3 (top) reports the average rank of each method and the corresponding statistical comparisons. The results confirm that PSDNorm significantly outperforms the baselines, underscoring the value of incorporating temporal structure into normalization for robust and data-efficient generalization. The same trend is observed for U-Sleep trained on all subjects (see in Appendix Figure˜8). The following experiments focus on the balanced@400 setup.
@@
PSDNorm is a plug-and-play normalization layer that can be seamlessly integrated into various neural network architectures. To demonstrate this flexibility, we evaluate its performance on both the U-Sleep and CNNTransformer models. Figure˜3 reports the average rank of each normalization method across datasets and subjects for both architectures using datasets balanced@400. In both architectures, PSDNorm achieves the best overall ranking and demonstrates statistically significant improvements over both BatchNorm, InstanceNorm, and TMA. The results confirm that PSDNorm generalizes well beyond a single architecture and can provide consistent improvements in diverse modeling setups which is not the case of TMA that is ranked the worst with CNNTransformer. InstanceNorm performs competitively in some cases but is never significantly better than PSDNorm. Detailed numerical scores for CNNTransformer are reported in the supplementary material (Table˜7).
It is important to highlight that PSDNorm brings improvements without too much additional computational cost. In appendix Section˜A.12 we provide a detailed comparison of the computational time of PSDNorm with other normalization layers. The results show that PSDNorm is only slightly slower than BatchNorm and InstanceNorm, with a negligible increase in training time (less than 10%) and no significant impact on inference speed.
@@
Performance variability across subjects is a key challenge in biomedical applications where ensuring consistently high performance—even for the most challenging subjects—is critical. To highlight the robustness of PSDNorm, Figure˜4 presents a scatter plot of subject-wise BACC scores comparing BatchNorm or InstanceNorm vs. PSDNorm across two selected target datasets. CHAT and MASS are two challenging datasets, where the prediction performance is significantly lower than the other datasets. For CHAT, most of the dots are below the diagonal, indicating that PSDNorm improves performance for 91% of subjects against BatchNorm and 99% of subjects against InstanceNorm, with the largest gains observed for the hardest subjects, reinforcing its ability to handle challenging cases. For MASS, PSDNorm improves performance for 75% of subjects against BatchNorm and 69% against InstanceNorm. This demonstrates that PSDNorm is not only effective in improving overall performance but also excels in enhancing the performance of the most challenging subjects.
4.3 Illustration of PSD Normalization
Figure 5 shows how different normalization layers affect the PSD of signals at several stages of the network. The input signals display limited variability, which explains why applying TMA as a pre-processing step provides only marginal benefit. In the first row, corresponding to BatchNorm, the PSD variability increases with depth, a behavior that is undesirable for generalization. TMA exhibits a similar pattern, as no normalization is applied within the network to counteract this accumulation of variance. In contrast, both InstanceNorm and PSDNorm reduce PSD variability across samples. However, InstanceNorm does not fully align the PSDs, and noticeable differences remain between samples. PSDNorm, on the other hand, achieves strong alignment of PSDs across samples, indicating its ability to normalize the underlying temporal correlations. This alignment is essential for improving robustness and generalization, particularly in settings involving distribution shifts.
Figure 5: Illustration of PSD normalization with different normalization layers. The figure shows the PSD of different segment of 17min from one subject batch as input, and after 3 encoders using different normalization layers.